OpenStudy (anonymous):

How do I go about solving: 1+Floor(Log2(x)) where x = 8

OpenStudy (kinggeorge):

First up, you need to find what \(\log_2(8)\) is. Can you tell me what that is?

OpenStudy (anonymous):

3 or 2^3?

OpenStudy (kinggeorge):

Right. It's 3. So now you have the equation \[1+\lfloor3\rfloor\]Can you finish it from here? Or do you need some help with the floor function?

OpenStudy (anonymous):

So it would be 4 then?

OpenStudy (kinggeorge):

Looks perfect to me.

OpenStudy (anonymous):

Okay so just to make sure I know what I'm doing. All you do is find what X would be and then add one to it?

OpenStudy (kinggeorge):

The floor function is a bit weird. For example, if you had \[1+\left\lfloor \log_2(9)\right\rfloor\]instead, you would simplify as follows. Since \(\log_2(9)\approx 3.17\), you have\[1+\left\lfloor \log_2(9)\right\rfloor\\ 1+\left\lfloor 3.17\right\rfloor \\ 1+3 \\ 4\] In general, the floor function gets rid of any numbers to the right of the decimal place.

OpenStudy (anonymous):

Ah alright. We're doing all this stuff by hand so my professor is making it pretty easy for us. Thank you KingGeorge, you've been a lot of help!

OpenStudy (kinggeorge):

Formally, \[\left\lfloor x\right\rfloor=\text{max}\{m\in\mathbb{Z} |\;m<x\]In laymans terms, the greatest integer that is not greater than \(x\).

OpenStudy (anonymous):

So following the same equation but making X 123,456 the answer would be 17.

OpenStudy (kinggeorge):

Looks perfect.

OpenStudy (anonymous):

Thanks again!

OpenStudy (kinggeorge):

You're welcome.